Anhydrous copper(II) sulfate is difficult to dry completely. What mass of copper(II) sulfate would remain after removing 90% of the water from 250. g of CuSO4×5H2O? 


1
Expert's answer
2022-03-13T15:19:20-0400

Molar mass=249•7

Molar mass of water in salt= 90

Mass of water in the salt= 90/249.7= 90.1g

(90/100)×90.1= 81.09= mass removed

Remaining mass= 250_-81.09= 168.9g